OpenStudy (anonymous):

A rocket travels for 1.35 seconds a distance of 19 meters. What is the rockets velocity?

OpenStudy (saifoo.khan):

\[Velocity = \frac{displacement}{time}\]

OpenStudy (anonymous):

what's displacement?

OpenStudy (saifoo.khan):

displacement = distance traveled in a specific direction.

OpenStudy (anonymous):

19 meters?

OpenStudy (saifoo.khan):

19/1.35

OpenStudy (anonymous):

thanks! & one more thinggg.

OpenStudy (anonymous):

wait, nvm. it's the same equation! :)

OpenStudy (saifoo.khan):

you welcome..

OpenStudy (anonymous):

velocity=speed+direction speed= distance/time VELOCITY= (19/1.35)+direction(examples: north, south,east, west, up, down..ect)

OpenStudy (anonymous):

BTW, 19/1.35= 14.07407407....